# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

### v1

@@ Appendix C Proofs in the smooth convex setting
In this section, we derive parameters that satisfy the sufficient conditions of Theorem B.4 in order to derive upper bounds for SGD in the convex setting. The propositions in this section can be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
C.1 Bounds for short step-sizes: $\gamma L<1$
@@
Proposition C.1 (Lyapunov parameters. Convex case, short step-sizes).
@@
  • $e_{t}=\frac{a_{t+1}\gamma^{2}\alpha_{t}}{\alpha_{t}-a_{t+1}\gamma^{2}}$,
@@
  • $\beta_{t}\equiv\beta=0$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4. Moreover, it holds that
@@
Proof.
Notice that the sufficient conditions from Theorem B.4 are homogeneous. Because our goal is to obtain bounds whose bias term is the smallest possible, we can without loss of generality impose that $a_{0}=1$, and try to maximize $\rho$. Through our numerical analysis, we empirically observe that we can take $\alpha_{t}\equiv\alpha>0$ and $\beta_{t}\equiv\beta=0$, thus justifying these choices. We dedicate the remainder of the proof to justify our remaining choices.
We start by focusing on Condition 5 from Theorem B.4, which reads
$\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq({\alpha}-a_{t+1}\gamma^{2})(a_{t}-a_{t+1}).$
Rearranging this inequality yields
@@
Introducing the temporary variables $u_{t}\coloneqq\tfrac{1}{L\alpha-\gamma a_{t}}$ and $c=\tfrac{\gamma}{(1-\gamma L)\alpha}$, we see that we obtained an arithmetic inequality $u_{t+1}\leq u_{t}-c$, such that $u_{t}\leq u_{0}-ct$. We know that $u_{0}=\tfrac{1}{L\alpha-\gamma}$, so we can write
@@
This choice respects the assumption $L\alpha>\gamma$ we made earlier, since $01-\gamma L>$ 0.
@@
$u_{t}$ $=$ $u_{0}-tc=\frac{TL}{1-\gamma L}-\frac{t\gamma}{\alpha(1-\gamma L)}=\frac{\alpha TL-t\gamma}{\alpha(1-\gamma L)}$
$=$ $\frac{\tfrac{1}{L}(1-\gamma L)+\gamma T-t\gamma}{\alpha(1-\gamma L)}=\gamma\frac{T-t-1+\tfrac{1}{\gamma L}}{\alpha(1-\gamma L)}.$
@@
To conclude the proof, we need to check all the sufficient Lyapunov conditions of Theorem B.4. Conditions 1, 2, 3 and 5 are already satisfied, by construction. It remains to check Conditions 4 and 6.
@@
thus concluding the proof. ∎
@@
Theorem C.2 (Bounds for SGD. Convex case, short step-sizes).
Let Assumptions A.4 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,1)$. Then for every $T\geq 1$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\gamma^{2}\sigma_{*}^{2}}{\rho(1-\gamma L)}\bar{e}_{T}^{\prime},$
where $\rho=2\gamma+\tfrac{2(1-\gamma L)}{LT}$ and
$\bar{e}_{T}^{\prime}=\frac{1}{T}\sum_{t=0}^{T-1}e_{t}^{\prime},\quad\text{with$e_{t}^{\prime}=\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}$}.$
Moreover, we have the simpler bounds
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{L\|x_{0}-x_{*}\|^{2}}{2\gamma LT+2(1-\gamma L)}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}.$
If moreover $f$ is convex, the above bounds hold true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.1 with Theorem B.4 and Lemma B.2. We moreover exploit the fact that $\beta_{t}=0$ in Proposition C.1, which allows us to replace Assumption A.5 with Assumption A.4 in Theorem B.4, see Remark B.5 for more details. In the statement, we introduced the notation $e_{t}^{\prime}$ such that $e_{t}=\tfrac{\gamma^{2}}{1-\gamma L}e_{t}^{\prime}$. ∎
Remark C.3.
We note that the parameters selected in Proposition C.1 include $\beta_{t}\equiv 0$, which, in line with Remark B.5, justify that we only require Assumption A.4 instead of Assumption A.5
C.2 Bounds for the optimal step-size: $\gamma L=1$
Considering the result of Theorem C.2 and letting $\gamma L\to 1$, we note that the bias term tends to $\tfrac{1}{2\gamma T}$ while the variance term diverges since its denominator is a multiple of $1-\gamma L$. This suggests there might be a complication to obtain tight bounds in the case $\gamma L=1$. As we will see later in Section F, we have empirical results showcasing that it is not possible to obtain a bound for SGD where the bias term is $\tfrac{1}{2\gamma T}$ for $\gamma L=1$ whilst preserving a finite variance term. Therefore, in this section, we will prove bounds where the bias term is given by $\frac{1}{\rho T}$, with $\rho<2\gamma$.
Proposition C.4 (Lyapunov parameters. Convex case, optimal step-size).
Let $\gamma L=1$ and $\varepsilon\in(0,2)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=(2-\varepsilon)\gamma$,
  • $a_{t}\equiv a=1$,
  • $\alpha_{t}=\gamma^{2}$,
  • $\beta_{t}=\tfrac{\gamma^{2}\varepsilon}{2}$,
  • $e_{t}=\gamma^{2}\tfrac{2+\varepsilon}{\varepsilon}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4.
Proof.
First, let us impose that $a_{0}=1$ without loss of generality. Second, we set $\rho=(2-\varepsilon)\gamma$ for $\varepsilon>0$ since numerical experiments suggest that $\rho=2\gamma$ is impossible (see Section F.2). Third, taking inspiration from our numerical findings, we impose that the sequence $a_{t}$ is constantly equal to $1$, that is $a_{t}\equiv a=1$.
If we want to satisfy the sufficient Lyapunov conditions of Theorem B.4, we need in particular to verify Condition 5, which reduces to
$\gamma=L\alpha_{t},$
because $a_{t}-a_{t+1}=0$. This means that $\alpha_{t}\equiv\alpha=\gamma^{2}$, since $\gamma L=1$. Condition 2 implies that $\beta_{t}$ must satisfy
$\beta_{t}\leq\alpha-\frac{\rho}{2L}=\frac{\gamma^{2}\varepsilon}{2}.$
Condition 4 is trivially satisfied since it requires
$\gamma^{2}\leq\alpha_{t}+\beta_{t}=\gamma^{2}+\beta_{t},$
that is $\beta_{t}\geq 0$. It remains to verify Condition 6, which requires that
$e_{t}\geq\frac{\gamma^{2}(\alpha+\beta_{t})}{\alpha+\beta_{t}-\gamma^{2}}.$
First, observe that this is the only constraint on $e_{t}$, which we want to see as small as possible. So we shall set the inequality to be an equality. Second, the expression is decreasing with respect to $\beta_{t}$ (its derivative has the same sign as $-\gamma^{2}$). Therefore, $e_{t}$ will be minimal if $\beta_{t}$ is maximal, meaning we shall fix $\beta_{t}$ to constantly be equal to its earlier derived upper bound, that is $\beta_{t}\equiv\beta=\tfrac{\gamma^{2}\varepsilon}{2}$. Finally, we compute
$e_{t}=\gamma^{2}\frac{\gamma^{2}+\tfrac{\gamma^{2}\varepsilon}{2}}{\gamma^{2}+\tfrac{\gamma^{2}\varepsilon}{2}-\gamma^{2}}=\gamma^{2}\frac{2+{\varepsilon}}{{\varepsilon}}.$
Note in particular that $e_{t}$ is constant here. ∎
Theorem C.5 (Bounds for SGD. Convex case, optimal step-size).
Let Assumptions A.5 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L=1$. Then, for every $T\geq 1$ and $\varepsilon\in(0,2)$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{(2-\varepsilon)\gamma T}+\frac{(2+\varepsilon)\gamma\sigma_{*}^{2}}{\varepsilon(2-\varepsilon)}.$
If moreover $f$ is convex, the above bound holds true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.4 with Theorem B.4 and Lemma B.2. ∎
@@ C.2 Bounds for the optimal step-size: $\gamma L=1$
In Section 4 we explained that our numerical experiments suggest that a bias with $\rho=2\gamma$, where $\gamma L=1$, is not achievable whilst keeping the variance finite. Theorem B.4 provides sufficient conditions on the parameters $\left(\rho,(a_{t}),(e_{t})\right)$ for the Lyapunov decrease to happen. Those sufficient conditions are the ones we use throughout the paper, and allow us to theoretically match the bounds obtained numerically. We therefore have good reasons to believe that those sufficient conditions are very close to be necessary. In what follows, we show that the system of inequalities in Theorem B.4 is not feasible for the optimal value $\rho=2\gamma$ when $\gamma L=1$.
Proposition C.6.
@@
This inequality is to be satisfied for all $\delta$ and $x_{t}$. Specifically, the inequality must hold for $(x_{t},\delta)=(1,0)$, showing that $a_{t+1}(1-\gamma L)^{2}-a_{t}+\rho L/2\leq 0$. Setting $\rho=2\gamma$, and $\gamma L=1$ yields $a_{t}\geq 1$. ∎
Corollary C.7.
The sufficient conditions of Theorem B.4 cannot be verified with $\rho=2\gamma$, $a_{0}=1$ and $\gamma L=1$.
Proof.
Let $\rho,(e_{t},a_{t})_{t=0}^{T}$ be parameters, normalized with $a_{0}=1$, and suppose $\gamma L=1$ and $\rho=2\gamma$. Suppose that there exists $(\alpha_{t},\beta_{t})$ such that the system of inequality in Theorem B.4 holds. From Proposition C.6, we know that $a_{t}\geq 1$ for all $t=0,\ldots,T$. Moreover, from Condition 3 and a simple induction argument, we deduce that $a_{t}\leq a_{0}=1$. Hence, $a_{t}\equiv 1$. This together with Condition 5 yields
@@
Finally by injecting the values of $\alpha_{t}$ and $\beta_{t}$ into Condition 6, we obtain
$\frac{1}{L^{4}}\leq 0.$
This contradicts $L>0$, thus concluding. ∎
C.3 Bounds for large step-sizes $\gamma L>1$
Proposition C.8 (Lyapunov parameters. Convex case, large step-sizes).
Let $21<\gamma L<$ 2, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
@@ C.3 Bounds for large step-sizes $\gamma L>1$
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4. Moreover, the averaged variance $\bar{e}_{T}\coloneqq\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ has the following asymptotic equivalence as $T\to+\infty$:
@@
Proof.
@@
$\rho=2L(\alpha_{t}+\beta_{t})\iff\alpha_{t}=\frac{\rho}{2L}+\beta_{t},$
@@
We observe an arithmetic-geometric relation given by, where we set $\theta=(\gamma L-1)^{2}$,
@@
We now verify the remaining conditions from Theorem B.4. The nonnegativity of $a_{t}$ implies the nonnegativity of $\alpha_{t},\beta_{t},$. Conditions 1, 3, 5 are verified by construction. It remains to check that Conditions 4 and 6 hold true.
@@
00 0 $\leq$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}-\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}$
@@
In particular, one sees that $\bar{e}_{T}^{\prime}\sim\frac{1}{T\theta^{T-1}}$. ∎

### v2

@@ A.1 Convex Setting
In this section, we derive parameters that satisfy the sufficient conditions of Theorem 3.1 in order to derive upper bounds for SGD in the convex setting. The propositions in this section can be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
A.1.1 Short Step-Sizes: $\gamma L<1$ (Proof of Theorem 4.1)
@@
Proposition 2(Lyapunov parameters. Convex case, short step-sizes)
@@
  • $\bullet$
@@
  • $\bullet$
    $\beta_{t}\equiv\beta=0$,
  • $\bullet$
    $e_{t}=\frac{a_{t+1}\gamma^{2}\alpha_{t}}{\alpha_{t}-a_{t+1}\gamma^{2}}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, it holds that
@@
In particular, the bound in Theorem 4.1 holds true.
Proof
Notice that the sufficient conditions from Theorem 3.1 are homogeneous. Because our goal is to obtain bounds whose bias term is the smallest possible, we can without loss of generality impose that $a_{0}=1$, and try to maximize $\rho$. Through our numerical analysis, we empirically observe that we can take $\alpha_{t}\equiv\alpha>0$ and $\beta_{t}\equiv\beta=0$, thus justifying these choices. We dedicate the remainder of the proof to justify our remaining choices.
We start by focusing on Condition 5 from Theorem 3.1, which reads
$\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq({\alpha}-a_{t+1}\gamma^{2})(a_{t}-a_{t+1})$
@@
Introducing the temporary variables $u_{t}\coloneqq\tfrac{1}{L\alpha-\gamma a_{t}}$ and $c=\tfrac{\gamma}{(1-\gamma L)\alpha}$, we see that we obtained an arithmetic inequality $u_{t+1}\leq u_{t}-c$, leading us to $u_{t}\leq u_{0}-ct$. We know that $u_{0}=\tfrac{1}{L\alpha-\gamma}$, so we can write
@@
This choice respects the assumption $L\alpha>\gamma$ we made earlier, since $1-\gamma L>0$.
@@
$u_{t}$ $=$ $u_{0}-tc=\frac{TL}{1-\gamma L}-\frac{t\gamma}{\alpha(1-\gamma L)}=\frac{\alpha TL-t\gamma}{\alpha(1-\gamma L)}=\frac{\tfrac{1}{L}(1-\gamma L)+\gamma T-t\gamma}{\alpha(1-\gamma L)}=\gamma\frac{T-t-1+\tfrac{1}{\gamma L}}{\alpha(1-\gamma L)}.$
@@
To conclude the proof, we need to check all the sufficient Lyapunov conditions of Theorem 3.1. Conditions 1, 2, 3 and 5 are already satisfied, by construction. It remains to check Conditions 4 and 6.
@@
thus concluding the proof. The bound in Theorem 4.1 follows by combining Theorem 3.1 and Lemma 7.
Note that the bound in Theorem 4.1 may be sharpened by using the exact values of $e_{t}$ rather than their bound. Finally, we prove the bias-optimality claim in Theorem 4.1.
@@
Proposition 3(Bias-optimality. Convex case, short step-sizes)
The bias obtained in the bound of Theorem 4.1 is optimal, in the sense that, for any $x_{0}\in\mathcal{H}$, there exists a convex and smooth problem for which $m=1$ and both inequalities are satisfied with equality. Moreover, this problem has bounded gradients.
Proof
We build our lower bound example in a similar to what was done in taylor_smooth_2017 . The idea is to consider Huber functions, which have minimizers while behaving linearly when far away from the minimizers. This linear behavior is not only slow, which fits our goal to find a worst-case function, but also allows for computing easily and explicitly the iterates. Given $\eta>0$, we define the function $\mathcal{H}_{\eta}$ as
$\mathcal{H}_{\eta}(x)=\begin{cases}\eta L\|x\|-\frac{L}{2}\eta^{2}&\text{if}\quad\|x\|>\eta\\ \frac{L}{2}\|x\|^{2}&\text{if}\quad\|x\|\leq\eta.\end{cases}$
This is exactly the classical Huber function multiplied by $L$, which is convex, $L$-smooth and has bounded gradients (bauschke_convex_2017, , Example 13.7). It now remains to prove the claim in item 3. As a preliminary, observe that if $\|x\|\geq\eta$ and if $x^{+}:=x-\gamma\nabla\mathcal{H}_{\eta}(x)$, then
[…]
Let $x_{0}\in\mathcal{H}$ and denote $\|x_{0}\|=D$, and let $(x_{t})$ be the iterates of gradient descent applied to $f=\mathcal{H}_{\eta}$, where the parameter $\eta$ is defined as $\eta=\tfrac{D}{1+(T-1)\gamma L}$. A simple induction argument shows that this choice of $\eta$ is small enough to ensure that $\|x_{t}\|\geq\eta$ for $t=0,\dots,T-1$, which in turn allows us to write $x_{t}=(1-t\tfrac{\gamma L\eta}{D})x_{0},$ and as such
$\overline{x}_{T}$ $=\frac{1}{T}\sum_{t=0}^{T-1}x_{t}=\frac{1}{T}\sum_{t=0}^{T-1}\left(1-t\gamma L\eta/D\right)x_{0}=\left(1-\frac{\gamma L(T-1)\eta}{2D}\right)x_{0}.$
A direct computation shows that $\|\bar{x}_{T}\|\leq\eta$, which means that
$f(\overline{x}_{T})-\inf f$ $=\eta L\left(D-\frac{\gamma L(T-1)\eta}{2}\right)-\frac{L}{2}\eta^{2}=\eta LD-L\eta^{2}\cdot\frac{1+\gamma L(T-1)}{2}=\eta LD-\frac{\eta LD}{2}=\frac{\eta LD}{2}=\frac{L\|x_{0}-x_{*}\|^{2}}{2(1+(T-1)\gamma L)},$
which indeed matches the bias term of Theorem 4.1. Moreover, the iterates $x_{0},\dots,x_{T}$ all belong to the segment $x_{0},x_{T}]$, on which $f$ is affine, which means that we also have $\frac{1}{T}\sum_{t=0}^{T-1}f(x_{t})-\inf f=f(\overline{x}_{T})-\inf f$, thus concluding.
A.1.2 Critical Step-Size: $\gamma L=1$
The proof of Theorem 4.2 is presented in Section A.1.4.
@@ A.1.2 Critical Step-Size: $\gamma L=1$
In Remark 4, we mentioned that we believe there to be a singularity at the critical step-size, making it such that the optimal bias cannot be achieved with a finite variance term. Moreover, this was confirmed numerically in Section 7. Theorem 3.1 provides sufficient conditions on the parameters $\left(\rho,(a_{t}),(e_{t})\right)$ for the Lyapunov decrease to happen. Those sufficient conditions are the ones we use throughout the paper, and allow us to theoretically match the bounds obtained numerically. We therefore have good reasons to believe that those sufficient conditions are very close to be necessary. In what follows, we show that the system of inequalities in Theorem 3.1 is not feasible for the optimal value $\rho=2\gamma$ when $\gamma L=1$.
Proposition 4
@@
This inequality is to be satisfied for all $\delta$ and $x_{t}$. Specifically, the inequality must hold for $(x_{t},\delta)=(1,0)$, showing that $a_{t+1}(1-\gamma L)^{2}-a_{t}+\rho L/2\leq 0$. Setting $\rho=2\gamma$, and $\gamma L=1$ yields $a_{t}\geq 1$.
Proposition 5
The sufficient conditions of Theorem 3.1 cannot be verified with $\rho=2\gamma$, $a_{0}=1$ and $\gamma L=1$.
Proof
Let $\rho,(e_{t},a_{t})_{t=0}^{T}$ be parameters, normalized with $a_{0}=1$, and suppose $\gamma L=1$ and $\rho=2\gamma$. Suppose that there exists $(\alpha_{t},\beta_{t})$ such that the system of inequality in Theorem 3.1 holds. From Proposition 4, we know that $a_{t}\geq 1$ for all $t=0,\ldots,T$. Moreover, from Condition 3 and a simple induction argument, we deduce that $a_{t}\leq a_{0}=1$. Hence, $a_{t}\equiv 1$. This together with Condition 5 yields
@@
Finally by injecting the values of $\alpha_{t}$ and $\beta_{t}$ into Condition 6, we obtain ${1}/{L^{4}}\leq 0,$ which contradicts $L>0$, thus concluding.
A.1.3 Large Step-Sizes: $\gamma L>1$ (Proof of Theorem 4.3)
Proposition 6(Lyapunov parameters. Convex case, large step-sizes)
Let $1<\gamma L<2$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
@@
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, the averaged variance $\bar{e}_{T}\coloneqq\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ has the following asymptotic equivalence as $T\to+\infty$:
@@
In particular, the bound in Theorem 4.3 holds true.
Proof
@@
$\rho=2L(\alpha_{t}-\beta_{t})\iff\alpha_{t}=\frac{\rho}{2L}+\beta_{t}.$
@@
We observe an arithmetic-geometric relation which can be summarized after setting $\theta=(\gamma L-1)^{2}$:
@@
We now verify the remaining conditions from Theorem 3.1. The nonnegativity of $a_{t}$ implies the nonnegativity of $\alpha_{t},\beta_{t}$. Conditions 1, 3, 5 are verified by construction. It remains to check that Conditions 4 and 6 hold true.
@@
$0$ $\leq$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}-\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}$
@@
In particular, one sees that $\bar{e}_{T}^{\prime}\sim\frac{1}{T\theta^{T-1}}$, where we recall that we defined $\theta=(\gamma L-1)^{2}\in(0,1)$. The bound in Theorem 4.3 now follows by combining Theorem 3.1 and Lemma 7.
We may moreover show the bias-optimality claim of Theorem 4.3 by providing a function that achieves the lower bound.